# Discovering universal temperature regulation dynamics in animals

### v1

@@ 5.1 Data sets
We used data sets from the following papers: [9, 14, 15, 16]. If the data was not available with the paper, we reached out to the authors directly.
@@ 5.3 A note on the spirit of these calculations
It should be noted that the dimension calculation is sensitive to changes in the data and the data-smoothing technique applied. It is also likely that the interpolation we performed on the black bear data had some impact on the dimension. However, our goal is not to definitively state the dimension of the effective thermophysiology system for any given species, but rather to get an understanding of coarse dimensionality (are the thermophysiological systems relatively low dimensional (2-5 dimensional) or relatively high dimensional (10-100 dimensional)?). Across species, we find that the dimensionality of thermophysiological systems is relatively low and can be qualitatively described using the models we lay out here, but that is not to say that alternative choices would not have slightly changed the dimensionality. Additionally, changing hyperparameters of MATLAB’s phaseSpaceReconstruction() would likely also change the dimensionality slightly. The ultimate test of dimensionality cannot be found in the statistics of an embedding algorithm or the associated hyperparameters of the algorithm. The ultimate test of dimensionality is whether the model can capture the observation. We were able to show this in a quantitative manner for the Arctic ground squirrel and qualitatively for all other species we examined.
@@ 5.4 Model selection
To perform model selection on Arctic ground squirrel body temperature data, we used the data found in 9903_M_Toolik_temp.csv in the supplement of [9]. We opted to use a segment of the Arctic ground squirrel body temperature data set that was six spikes long. Specifically, we used from index 10811 to index 13747 as training data and from index 13748 to index 16573 as validation data. We applied min-max scaling to the training and validation data. We note that using different data sets may produce different model selection results (see Discussion for further information on modeling choices and limitations).
@@ 5.5 Identifiability
To ensure that the entire suite of models defined by Eqs. (1)-(2) were structurally identifiable ahead of parameter estimation, we used StructuralIdentifiability.jl [35] to check that Model 10 was identifiable after setting $\theta_{4}$ to its numerical value found during model selection (specifically the value of $\theta_{4}$ for the eight-term model recovered during model selection, though the values of $\theta_{4}$ in the nine-term and ten-term model are almost identical). We found that all parameters were globally identifiable with $\theta_{4}$ fixed. If Model 10 is identifiable, then our other models of interest, which are subsets of Model 10, will be as well. It should be noted that there is no identifiability guarantee for models we found during model selection [36].
@@ 6.1 Figures
Figure 5: Plane plane for Equations (5)-(6). The $\dot{\hat{x}}=0$ nullcline is shown in blue, and the $\dot{\hat{y}}=0$ is shown in orange. The limit cycle is shown in pink for $\gamma=1.2$, $\kappa=0.02$, $\iota=2$ x 10-6.
@@
Figure 8: Cost function value (also known as “action”) vs. $\beta$ plots at different sparsity thresholds (purple). An action vs. $\beta$ curve that is plateaued indicates an appropriate balance of data error and model error enforcement. For this reason, we examined models for $75\leq\beta\leq 85$ in our study. These cutoffs are delineated by the vertical yellow lines.
@@ 6.3 Nondimensionalization and assumptions on model parameters
where $\gamma=\dfrac{\sqrt{\dfrac{\theta_{5}}{\theta_{6}}}\theta_{3}}{\theta_{1}}$, $\xi=\dfrac{\sqrt{\dfrac{\theta_{5}}{\theta_{6}}}\theta_{4}}{\theta_{1}}$, $\iota=\dfrac{\theta_{0}(\theta_{2})^{2}}{(\theta_{1})^{3}}$, and $\kappa=\dfrac{\sqrt{\theta_{5}\theta_{6}}}{\theta_{1}}$.
[…]
@@
6.4 Modified model functional forms
@@ 6.4.3 The black bear model
C. E. F. is supported in part by The James S. McDonnell Foundation Postdoctoral Fellowship Award in Complex Systems (https://doi.org/10.37717/2020-1591) and by the NSF-Simons Center for Quantitative Biology at Northwestern University (NSF: 1764421 and Simons Foundation/SFARI 597491-RWC). N. M. M. was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DESC0024253. A.J.E. was supported as a Quantitative Biology REU student at Northwestern University (NSF: 2150134). This research was supported in part by grants from the NSF (DMS-2235451) and Simons Foundation (MP-TMPS-00005320) to the NSF-Simons National Institute for Theory and Mathematics in Biology (NITMB).
@@ Acknowledgments
We thank Mary Silber, Danny Abrams, and Bard Ermentrout for helpful discussions regarding bifurcation theory and dynamical systems, Roan Luikart for his help digitizing data sets, and Alasdair Hastewell for his careful read and suggestions on the manuscript. C.E.F. also thanks the organizers and participants of the 2023 University of Alaska Fairbanks Hibernation Science Workshop.

### v2

@@ 6.1 Data sets
We used data sets from the following papers: [9, 21, 22, 23]. If the data were not available with the paper, we reached out to the authors directly.
@@ 6.3 A note on the spirit of these calculations
It should be noted that the dimension calculation is sensitive to changes in the data and the data-smoothing technique applied. It is also likely that the interpolation we performed on the black bear data had some impact on the dimension. However, our goal is not to definitively state the dimension of the effective thermophysiology system for any given species, but rather to get an understanding of coarse dimensionality (are the thermophysiological systems relatively low-dimensional (2-5 dimensional) or relatively high-dimensional (10-100 dimensional)?). Across species, we find that the dimensionality of thermophysiological systems is relatively low and can be qualitatively described using the models we lay out here, but that is not to say that alternative choices would not have slightly changed the dimensionality. Additionally, changing hyperparameters of MATLAB’s phaseSpaceReconstruction() would likely also change the dimensionality slightly. The ultimate test of dimensionality cannot be found in the statistics of an embedding algorithm or the associated hyperparameters of the algorithm. The ultimate test of dimensionality is whether the model can capture the observation. We were able to show this in a quantitative manner for the Arctic ground squirrel and qualitatively for all other species we examined.
@@ 6.4 Model selection
To perform model selection on Arctic ground squirrel body temperature data, we used the data found in 9903_M_Toolik_temp.csv in the supplement of [9]. We opted to use a segment of the Arctic ground squirrel body temperature data set that was six spikes long. Specifically, we used the data from index 10811 to index 13747 as training data and from index 13748 to index 16573 as validation data. We applied min-max scaling to the training and validation data. We note that using different data sets may produce different model selection results (see Discussion for further information on modeling choices and limitations).
@@ 6.5 Identifiability
To ensure that the entire suite of models defined by Eqs. (1)-(2) were structurally identifiable ahead of parameter estimation, we used StructuralIdentifiability.jl [43] to check that Model 10 was identifiable after setting $\theta_{4}$ to its numerical value found during model selection (specifically the value of $\theta_{4}$ for the eight-term model recovered during model selection, though the values of $\theta_{4}$ in the nine-term and ten-term models are almost identical). We found that all parameters were globally identifiable with $\theta_{4}$ fixed. If Model 10 is identifiable, then our other models of interest, which are subsets of Model 10, will be as well. It should be noted that there is no identifiability guarantee for the models we found during model selection [44].
@@ 7.1 Figures
Figure 8: Cost function value (also known as “action”) vs. $\beta$ plots at different sparsity thresholds (purple). An action vs. $\beta$ curve that is plateaued indicates an appropriate balance of data error and model error enforcement. For this reason, we examined models for $75\leq\beta\leq 85$ in our study (cutoffs shown as yellow lines).
@@ 7.3 Nondimensionalization and assumptions on model parameters
where $\gamma=\dfrac{\sqrt{\dfrac{\theta_{5}}{\theta_{6}}}\theta_{3}}{\theta_{1}}$, $\xi=\dfrac{\sqrt{\dfrac{\theta_{5}}{\theta_{6}}}\theta_{4}}{\theta_{2}}$, $\iota=\dfrac{\theta_{0}(\theta_{2})^{2}}{(\theta_{1})^{3}}$, and $\kappa=\dfrac{\sqrt{\theta_{5}\theta_{6}}}{\theta_{1}}$.
We make the assumption $\gamma\approx\xi$. During parameter estimation, we fix $\theta_{4}$ to make the models identifiable ($\theta_{4}=9.99995$, the value of Model 8AB that came out of model selection with $\theta_{0}=0$). After reoptimization, the numerical values of the nondimensionalized parameters are $\gamma=0.9930155091$, $\xi=0.9776338555$, $\iota=0.00002991498959$, and $\kappa=0.2621331463$. From this, we can see that $\gamma\approx\xi$. Under this assumption, we can factor the system, which yields the model we analyze:
@@
In Section 3.5, we used the following scalings $\hat{x}=\dfrac{x}{s_{1}}$, $\hat{y}=\dfrac{y}{s_{2}}$, and $\tau=\dfrac{t}{s_{0}}$ with $s_{1}=\dfrac{\theta_{3}}{\theta_{4}}$, $s_{2}=\dfrac{\sqrt{\theta_{5}\theta_{6}}\theta_{3}}{\theta_{6}\theta_{4}}$, and $s_{0}=\dfrac{(\theta_{4})^{2}}{(\theta_{3})^{2}\theta_{2}}$ to find the dimensionless system described by Eqs. (5)-(6).
7.4 Modified model functional forms and parameter values
The Arctic ground squirrel model
In Figure 5D, we simulate Eqs. (10)-(11) using the parameter values $\gamma=1.1,$ $\kappa=0.1,$ $\iota=1$ x 10-4, $\hat{x}_{0}=1$ x 10-4, and $\hat{y}_{0}=1$ x 10-4.
@@ 7.4.3 The black bear model
C. E. F. is supported in part by The James S. McDonnell Foundation Postdoctoral Fellowship Award in Complex Systems (https://doi.org/10.37717/2020-1591) and by the NSF-Simons Center for Quantitative Biology at Northwestern University (NSF: 1764421 and Simons Foundation/SFARI 597491-RWC). N. M. M. was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, under Award Number DESC0024253. A.J.E. was supported as a Quantitative Biology REU student at Northwestern University (NSF: 2150134). This research was supported in part by grants from the NSF (DMS-2235451) and Simons Foundation (MP-TMPS-00005320) to the NSF-Simons National Institute for Theory and Mathematics in Biology (NITMB). This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
@@ Acknowledgments
We thank Mary Silber, Danny Abrams, and Bard Ermentrout for helpful discussions regarding bifurcation theory and dynamical systems, Roan Luikart for his help digitizing data sets, and Alasdair Hastewell for his careful read and suggestions on the manuscript. C.E.F. also thanks the organizers and participants of the 2023 University of Alaska Fairbanks Hibernation Science Workshop. C.E.F. thanks Matthew Gorby of Research Computing Services at Northwestern University for computational assistance.